A battery powered toy car accelerates from 0.0 m/s to 8.5 m/s. It’s mass is 0.55 kg. The chemical store in its battery is decreased by 30 J. Calculate the efficiency of this process.


1
Expert's answer
2021-05-23T16:35:05-0400
"\eta=\frac{mv^2}{2E}\\\eta=\frac{(0.55)(8.5)^2}{2(30)}=0.66"